OpenStudy (anonymous):

Determine the equation of the line, in slope-intercept form, that will get your spacecraft from Point B to Point C. Point B: (-3, 0), Point C: (-1, -4)........HELP PLEASE

OpenStudy (imstuck):

First find the slope using the formula\[\frac{ y _{2} -y _{1}}{ x _{2}-x _{1} }\]That would be \[\frac{ -4-0 }{ -1-(-3) }=\frac{ -4 }{ 2 }=-2\]So the slope is -2. The slope-intercept form of an equation is y = mx + b, where m is the slope and b is the y intercept. But in order to get there we have to use the point-slope formula because we have points and no y intercept). The point-slope formula is\[y-y _{1}=m(x-x _{1})\]where m is the slope, y1 is your y coordinate and x1 is your x coordinate of the same given point. Let's fill this in now using the point (-3, 0).\[y-0=-2(x-(-3))\]Doing the math you get y - 0 = -2x - 6. Now add 0 to both sides giving you y = -2x - 6. And that's your answer in slope-intercept form! The slope is -2, and the y intercept is -6.

OpenStudy (anonymous):

omg thank you so much you made it sounds so much easier than my online school........THANKS AGAIN

OpenStudy (imstuck):

You're so very welcome! I love helping (that's why I'm a teacher!) : )